OpenStudy (anonymous):

Choose the correct simplification of (8x4y3)2. 16x6y5 64x8y6 64x6y5 16x8y6

OpenStudy (anonymous):

PLZ CHECK THE QUESTION

OpenStudy (phi):

\[ (8x^4y^3)^2 = 8x^4y^3\cdot 8x^4y^3\]

OpenStudy (ajprincess):

\((8x4y3)^2=8^2*(x^4)^2*(y^3)^2\) \(=64*(x^{4*2})*(y^{3*2})\)

OpenStudy (anonymous):

IF YOUR QUESTION IS AS @phi did then HERE is the HINT: